OpenStudy (anonymous):

Let f be the function given by f (x) =3x+|4-x| Find and simplify f(5)+[f(-1)]^2

OpenStudy (amoodarya):

f (x) =3x+|4-x| put x=5 in it then put x=-1 then f(5)+(f(-1))^2

OpenStudy (amoodarya):

hint : f(5)=3*5+|4-5|=15+1=16 now can you go on?

OpenStudy (anonymous):

For this question they gave me a answer key. The answer for this question is 20. Can you show me how to get this answer please.

jigglypuff314 (jigglypuff314):

same concept as amoodarya was getting at :) f(-1) = 3(-1) + |4 - (-1)| = -3 + | 5 | = 2 then plug in what you know f(5) + ( f(-1) )^2 = 16 + (2)^2 = 16 + 4 =

OpenStudy (anonymous):

Thank you for your help, Jigglypuff314. Your answer was very clear.